A roller coaster car is going over the top of a 16-m-radius circular rise. At the top of the hill, the passengers "feel light," with an apparent weight only 60% of their true weight. How fast is the coaster moving?
Answer #1

at the highest point due to his apparent weight , \

net weight = 100-60   = 40 % real weight = 0.4

now use a = g = V^2/r

V^2 = 0.4* 16 *9.8

V = 7.919 m/s